OpenStudy (anonymous):

6x+8y=6 x-32y=-99 Solve using elimination, Plese Help

OpenStudy (anonymous):

multiply the first equation by \(4\) and start with \[24x+32y=24\\ x-32y=-99\] then add them up

OpenStudy (anonymous):

Why 4?

OpenStudy (anonymous):

add them up and you will see why 4

OpenStudy (anonymous):

Ok I asked so I know how to start other problems. how do you know what number to multiply by

OpenStudy (anonymous):

do you see that when you add the \(y\) terms go away? that was the point

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[24x+32y=24\\ x-32y=-99\] adding gives \[25x=-75\] no y's so you can solve for \(x\) alone

OpenStudy (anonymous):

multiplying by 4 is what made that work you have to figure that part out on your ownb what number would you multiply by to make one variable go away when you add

OpenStudy (anonymous):

sometimes you have to multiply both equations by some number not in this case because 32 is a multiple of 8

OpenStudy (anonymous):

oohh ok. so i got -3x and to solve for y I do the same?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

first off i assume you mean \(x=-3\) right?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

then since you know what x is , you do not have to redo the procedure replace x by -3 in either equation and solve for y

OpenStudy (anonymous):

\[x-32y=-99\\ -3-32x=-99\] etc

OpenStudy (anonymous):

(-3,3)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks A lot